OpenStudy (anonymous):

Show that {an} is monotonic and determine whether {an} is convergent or divergent. If the sequence is convergent, find its limit.

OpenStudy (anonymous):

\[a) \left\{ a _{n} \right\}=\left\{ \frac{ \sqrt{n} }{ 1+\sqrt{n} } \right\}\]

OpenStudy (anonymous):

\[b) \left\{ a _{n} \right\} = \left\{ 4+ \frac{ 1 }{ 2^{n} } \right\}\]

OpenStudy (experimentx):

post your other question separetely

OpenStudy (experimentx):

define a function f(x) = sqrt(x)/(1 + sqrt(x)) do second derivative test. or induction might be helpful to show the monotocity.

OpenStudy (anonymous):

how do you show the monotocity?

OpenStudy (experimentx):

how do you show that function is increasing?

OpenStudy (anonymous):

so i would have to use the quotient rule to determine the derivative of f(x) rigt?

OpenStudy (experimentx):

you can show that the slope of the function is increasing.

OpenStudy (experimentx):

or you can show that via induction.

OpenStudy (anonymous):

i would use the derivative of f(x) to show if its increasing or decreasing.

OpenStudy (experimentx):

yes you can do that ... show that the slope is greater than zero for all x>1 for \[f(x) = \frac{\sqrt{x}}{1 + \sqrt{x}}\]

OpenStudy (anonymous):

so when i have my derivative, how would i be able to tell that the slope is greater for zero for all x>1?

OpenStudy (anonymous):

well i know the denominator will always be positive because it's squared

OpenStudy (anonymous):

and the tops have square roots in them.

OpenStudy (anonymous):

is my sequence bounded? how do i tell?

OpenStudy (experimentx):

just find a suitable upper bound for your sequence

OpenStudy (anonymous):

i dont understand.

OpenStudy (experimentx):

you can show that this sequence is always less than one.